Citation Builders, Inc., builds office buildings and single-family homes. The office buildings are constructed under contract with reputable buyers. The homes are constructed in developments ranging from 10-20 homes and are typically sold during construction or soon after. To secure the home upon completion, buyers must pay a deposit of 10% of the price of the home with the remaining balance due upon completion of the house and transfer of title. Failure to pay the full amount results in forfeiture of the down payment. Occasionally, homes remain unsold for as long as three months after construction. In these situations, sales price reductions are used to promote the sale. During 2018, Citation began construction of an office building for Altamont Corporation. The total contract price is $20 million. Costs incurred, estimated costs to complete at year-end, billings, and cash collections for the life of the contract are as follows: 2018 $ 4,000,000 12,000,000 2,000,000 1,800,000 2019 2020 9,500,000 4,500,000 10,000,000 8,600,000 Costs incurred during the year Estimated costs to complete as of year-end Billings during the year Cash collections during the year $4,500,000 8,000,000 9,600,000 Also during 2018, Citation began a development consisting of 12 identical homes. Citation estimated that each home will sell for $600,000, but individual sales prices are negotiated with buyers. Deposits were received for eight of the homes, three of which were completed during 2018 and paid for in full for $600,000 each by the buyers. The completed homes cost $450,000 each to construct The construction costs incurred during 2018 for the nine uncompleted homes totaled $2,700,000. Required: 1. Which method is most equivalent to recognizing revenue at the point of delivery? 2. Answer the following questions assuming that Citation uses the completed contract 2-a. How much revenue related to this contract will Citation report in its 2018 and 2019 income statements? 2-b. What is the amount of gross profit or loss to be recognized for the Altamont contract during 2018 and 2019? 2-c. What will Citation report in its December 31, 2018, balance sheet related to this contract? (lgnore cash.) 3. Answer the following questions assuming that Citation uses the percentage-of-completion method for its office building contracts. 3-a. How much revenue related to this contract will Citation report in its 2018 and 2019 income statements? 3-b. What is the amount of gross profit or loss to be recognized for the Altamont contract during 2018 and 2019? for its office building contracts:
Complete this question by entering your answers in the tabs below. Req 1 2A and 2B Req 3A and Req 4A and 4B Req 2C Req 3C Req 5 Req 4C Req 6A Req 6B 3B Assume the same information for 2018 and 2019, but that as of year-end 2019 the estimated cost to complete the office building is $9,000,000. Citation uses the percentage-of-completion method for its office building contracts. How much revenue related to this contract will Citation report in the 2019 income statement? What is the amount of gross profit or loss to be recognized for the Altamont contract during 2019? (Loss amounts should be indicated with a minus sign.) Show less A 2019 $7,000,000 Revenue reported $ 2,000,000 (x Gross profit or loss to be recognized

Req 1 2A and Req 3A and 3B Req 4A and Req 2C Req 3C Req 4C Req 5 Req 6B Req 6A 2B 4B What will Citation report in its 2018 income statement related to the single-family home business (ignore cash in the balance sheet)? Income Statement (Partial) 2018 Loss on long-term contract Loss on long-term contract Gross profit $ 0
Balance Sheet (Partial) At December 31, 2018 Current assets: Accounts receivable Current liabilities: Billings on construction contract
Answer #1

Part 4 A & B

Revenue reported

$7000000

Gross profit or loss to be recognized

$(3500000)

2018

2019

Contract price

20000000

20000000

Actual costs to date

4000000

13500000

(9500000+4000000)

Estimated costs to complete

12000000

9000000

Total estimated costs

16000000

22500000

Estimated gross profit

4000000

(2500000)

A. Revenue recognition:

Total revenue recognized to date= (percentage complete) ×(total revenue)

= (13500000/22500000) x 20000000 = 12000000

Revenue recognized in 2019= total – revenue recognized in prior periods = 12000000-5000000 = $7000000

(Revenue recognized in prior period (2018) = 4000000/16000000*20000000 = 5000000)

B. Gross profit or loss recognition:

Gross loss = overall loss – previously recognized gross profit = (2500000)-1000000 = (3500000)

(previously recognized gross profit (2018) = 5000000-4000000=1000000)

Part 4 C

Current assets:

Accounts receivable

1600000

Current liabilities:

Billings in excess of Costs and profit

1000000

Accounts receivable = (2000000-1800000)+(10000000-8600000) = 1600000

Citation will report in its December 31, 2019, balance sheet related to this contract = billings – (cost + profit ) = (2000000+10000000)-(4000000+9500000+1000000-3500000) = 1000000

Part 6 A

Income Statement (Partial)

Sales revenue (600000*3)

1800000

Cost of goods sold (450000*3)

1350000

Gross Profit

$450000

Part 6 B

Balance Sheet (Partial)

Current assets:

Inventory (work in process)

2700000

Current liabilities:

Customer deposits (or unearned revenue)

300000

Customer deposits = 600000*10%*5